OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y = ±three divided by four times x.

OpenStudy (anonymous):

I got \[\Large \frac {y^2}{36}-\frac {x^2}{64}=1\]

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@precal

OpenStudy (dumbcow):

yes you are correct

OpenStudy (anonymous):

Thnks

hero (hero):

@doulikepiecauseidont , if you want to check your work all you have to do is use this site: https://www.desmos.com/calculator/j5fdwo0xv0

OpenStudy (anonymous):

Thanks! @Hero